. 1. A torsion pendulum consists of a 2.0 Kg solid disk of 0.15m radius. When a 1 N.m torque is applied it displaces 150. What is its frequency of vibration? The moment of inertia for a disk is: Conversion:
Answer #1

Moment of inertia = 1/2mr^2= 1/2× 2× 0.15×0.15= 0.0225m^2

Torque = torsional constant X angle of twist

Torsional constant = a

Angle of twist = 15degree= 0.2617radians

1= a X 0.2617

a= 1/0.2617= 3.82Nm/radian

T = 2×pi ( I/ a)^1/2

T = 2×3.14 (0.0225/3.82)^1/2

T= 0.0482s

f=20.7Hz